The length (๐‘™) and Time Period (๐‘‡) of a simple pendulum are related as: ๐‘‡ = 2๐œ‹โˆš

๐‘™

๐‘”

, whereย 

๐‘” is acceleration due to gravity. In an experiment the following values for ๐‘™ and ๐‘‡ are found:

๐‘™ (in cm) ๐‘‡ (in seconds) ๐‘™ (in cm) ๐‘‡ (in seconds) ๐‘™ (in cm) ๐‘‡ (in seconds)

50 1.41 70 1.67 90 1.90

55 1.48 75 1.73 95 1.95

60 1.55 80 1.79 100 2.00

65 1.61 85 1.84

You can verify, if you like, taking ๐‘” โ‰ˆ 981 ๐‘๐‘š ๐‘ 

โˆ’2

. The values of ๐‘‡ are approximated to theย 

second place after decimal. Draw the ๐‘™ โˆ’ ๐‘‡ graph (do it manually, donโ€™t use a computer


1
Expert's answer
2021-03-29T09:03:04-0400

Explanations & Calculations


  • This experiment is done to verify the value of gravitational acceleration: g & the data is usually plotted as "\small T^2 \, \text{vs}\, L".
  • And By the gradient of the graph, the value of g can be calculated.
  • To make the hand drawing of the plot easier, the values of "\small T^2" can be written in whole numbers & powers as done here.


  • Refer to the figures attached.


  • Table of data



  • Plot



  • Omit any extraneous values that may be some mistakes (such as L:80cm & T:1.70s) & plot the graph.


  • To calculate the gradient of the graph, try to find farthest 2 points that lie on the best fit line & calculate the tangent value.

"\qquad\qquad
\begin{aligned}
\small m&=\small \tan \theta =\frac{(400-219)\times10^{-2 }}{1-0.55}\\
&=\small 402.22\times 10^{-2}
\end{aligned}"

  • Then compare the values,

"\qquad\qquad
\begin{aligned}
\small T&=\small 2\pi \sqrt{\frac{L+\epsilon}{g}}\\
\small T^2 &=\small 4\pi^2 \frac{(L+\epsilon)}{g}\\
\small T^2 &=\small \frac{4\pi^2}{g}\cdot L+\frac{4\pi^2 }{g}\cdot\epsilon\\
&\downarrow\\
\small y&=\small mx+c\\
\\
 \small m&=\small 4.0222=\frac{4\pi^2}{g}\\
\small g&=\small \frac{4\pi^2}{4.0222}=9.815\\
&\approx\small \bold{9.82ms^{-2}}
\end{aligned}" "\small L+\epsilon :" is the actual distance to the center of the pendulum


  • Experiment if satisfactorily performed.